OpenStudy (help_people):

Given the following functions f(x) and g(x), solve f[g(6)] and select the correct answer below: f(x) = 6x + 12 g(x) = x − 8 −96 0 24 48

OpenStudy (help_people):

@mathstudent55

OpenStudy (mathstudent55):

Find g(6) first using function g(x)

OpenStudy (mathstudent55):

g(x) = x - 8 g(6) = 6 - 8 What is 6 - 8?

OpenStudy (mathstudent55):

@help_people Are you there?

OpenStudy (lynfran):

solve for g(6) first g(x)=x-8 g(6)=6-8 =-2 f(x)=6x+12 f(gx)= 6(-2)+12 =-12+12 =0

OpenStudy (mathstudent55):

Once you find the value of g(6), use that value in function f(x).

OpenStudy (help_people):

-2 srry my mom called me

OpenStudy (help_people):

@mathstudent55

OpenStudy (mathstudent55):

Great. You are correct. g(6) = -2 Now we take -2 and use it in function f f(x) = 6x + 12 f(-2) = 6(-2) + 12 Now evaluate the line above.

OpenStudy (help_people):

ok

OpenStudy (help_people):

f(-2)=-12+12 f(-2)=0 is that the answer?

OpenStudy (help_people):

@mathstudent55

OpenStudy (mathstudent55):

Exactly. Good job!

OpenStudy (help_people):

Thank you! XOXOXO thanks for everything :)

OpenStudy (mathstudent55):

You're welcome.